A Bibliography of Papers in Lecture Notes in Computer Science (2006)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.09

Title word cross-reference

2 [55]. 2t + 1 [53]. p [38]. -Adic [55]. -Polynomials [38]. 0 [49]. 1 [48, 28]. 10th [167, 107, 154, 173]. 11th [150, 229, 140]. 12th [85, 190, 174, 231, 155]. 13th [105, 179]. 15th [159, 177, 68, 178]. 16th [183, 97, 118]. 18th [201]. 19th [152]. 2005 [205]. 2006 [115, 154]. 23rd [142]. 25th [94]. 28th [187]. 3-Designs [20]. 32nd [100]. 3rd [182].

4-7 [230]. 4th [70, 184, 206, 121, 204, 211, 108, 123, 132]. 5th [220, 209, 166, 73, 69, 217, 84, 81]. 6/WG [201]. 6th [215, 111, 90, 218, 202, 222, 112, 139, 164, 88, 82, 228, 147, 86, 158, 168]. 7th [227, 76, 130, 145, 181, 77, 136, 117, 113, 114, 79, 129, 199]. 8.8/11.2 [181]. 8th [78, 225, 224, 193, 93, 182, 119, 186, 103]. 9th [175, 151, 176, 163, 180, 124, 197]. AAECC [118]. AAECC-16 [118].

Abstract [117]. Access [31]. ACCV [113, 114]. ACSAC [229]. Action [182]. Ad [87]. Ad-hoc [87]. Adaption [156]. Adaptive [135]. Addition [52]. Additional [59]. ADG [84]. Adic [55]. Advanced [181, 115, 140, 202, 104]. advances [119, 89, 195, 154, 229, 231, 187, 173, 228, 207, 101]. Affine [33]. Against [64]. Agent [162, 122, 169, 149, 148, 158, 156, 99]. Agile [227]. AI [214]. Algebra [118, 182]. Algebraic [24, 118, 53]. Algorithm [18, 10, 1]. Algorithms [137, 92, 118, 174, 63, 220, 71, 218, 97, 219]. Allocation [9, 219]. Almost [41]. American [145]. AMR [135]. Amsterdam [96, 78]. Analyses [92]. Analysis [66, 174, 53, 132, 82, 147, 13, 12, 196, 130, 75]. Angeles [74]. Annual [167, 142]. Antipolis [107]. AOM [106]. Appendix [13, 12]. Application [146, 55, 231, 32, 107]. Applications [111, 198, 48, 181, 222, 69, 203, 21, 65, 140, 171, 182, 131, 1, 221, 148, 71, 74, 228, 104, 86, 169, 165, 83]. Applied [168, 77, 118]. Approaches [176, 217, 115, 213]. approximation [71, 137]. April [167, 144, 172, 161, 185, 163, 181, 92, 164, 193, 174, 187, 140, 171, 170, 173, 191, 165, 143, 179, 197]. APWeb [104, 103]. Architecture [146, 229, 152]. Architectures [141]. ARCS [152]. Areas [155]. Arrays [52]. Art [77]. Artificial [183, 143, 129]. artificielle [129]. Asia [229, 173, 103]. Asia-Pacific [229, 103]. ASIACRYPT [231]. Asian [113, 114]. aspect [138]. aspect-oriented [138]. Aspects [125, 142, 93]. Assignment [31]. Athens [222, 116]. Attack [51]. Attacking [65]. Attacks [24, 50, 28, 29]. Attention [157]. August [73, 146, 155, 194, 86]. Austria [175, 176, 112, 174, 210, 177, 178, 179]. Autocorrelation [43]. automata [107]. automated [73, 84]. Autonomic [116]. Aveiro [82]. Awareness [212, 62].

B [13]. Back [14]. Bangalore [88]. Barbara [180]. Barcelona [76, 92]. barriers [213]. Based [110, 38, 59, 224, 21, 159, 64, 52, 60, 35, 37, 110, 149]. Bases [56]. Bay [106]. Beijing [198, 95]. Being [200]. Berder [139]. Bergen [205]. Berlekamp [18]. Bern [206]. Bertinoro [202]. Between [36, 62]. Bilinear [37]. Binary [19, 21, 15, 43]. BIOADIT [115]. BioDM [171]. Biological [82]. Biologically [115]. Biology [167, 188]. Biomedical [171]. Biometric [54]. biometrics [101]. Blind [147]. Bohinj [189]. Bologna [225, 224]. Boolean [25, 53]. Bounds [46]. BPD [91]. BPI [91]. BPM [91]. BPRM [91]. BPS [91]. Braid [30]. Branches [33]. Brazil [199]. Bremen [123]. Broadcast [58]. Budapest [163, 121, 164, 165]. Bulgaria [81]. Business [91].

CA [74, 223, 180, 208, 120]. Cala [220]. Canada [150, 182, 155]. Card [181]. CARDIS [181]. CASSIS [196]. Castle [109]. Catharines [182]. CC [177]. CG [108]. Chains [32]. Chapter [7, 8, 10, 4, 5, 9]. Characteristic [25]. Characteristics [48]. Charleston [93, 117, 147]. Charting [131]. Checking [117, 179]. Chengdu [207]. Chile [145]. China [198, 161, 97, 98, 95, 229, 87, 231, 191, 104, 207, 101, 103]. Chosen [64]. Chosen-Ciphertext [64]. CIAA [107]. CIAC [218]. CICLing [136]. CiE [213]. Ciphers [24, 52]. Ciphertext [64]. Ciphertexts [58]. CISC [95]. City [136]. Class [36]. Classes [43]. classical [71]. CLIMA [158]. CM [55]. CMMR [160]. Code [21]. Codes [19, 118, 20, 18, 22, 23, 36, 16, 17, 15]. Coding [205]. Coefficients [5]. coherence [75]. Coimbra [209]. College [68]. Collision [49]. Collisions [50]. Combinatorial [214, 164]. combinatorical [71]. Communicating [201].

Communication [209, 202, 216, 116]. Communications [209, 206, 9]. Community [70, 166]. Compiler [177]. compilers [68]. Complementary [40]. Complete [31]. Complexities [44]. Complexity [218]. Component [147]. Computability [213]. Computation [175, 198, 164, 180, 97]. Computational [80, 167, 188, 213, 215, 136, 158, 199]. Computer [209, 202, 142, 139, 153, 203, 210, 182, 113, 114, 94, 100, 229, 203, 210, 160, 113, 114, 124]. Computers [108]. Computing [195, 185, 77, 204, 112, 211, 223, 152, 23, 79, 165, 85, 89, 96, 150, 81, 230, 68]. Concept [132]. Concepts [13, 12]. Conditions [22]. Conference [227, 175, 80, 215, 167, 176, 214, 213, 209, 206, 106, 198, 218, 161, 195, 225, 185, 163, 181, 222, 92, 95, 204, 211, 136, 223, 224, 164, 152, 134, 108, 174, 200, 154, 87, 231, 187, 140, 88, 208, 132, 216, 177, 117, 113, 114, 173, 191, 228, 120, 94, 147, 124, 143, 133, 201, 100, 186, 168, 197, 103, 85, 96, 78, 73, 74, 107, 229, 123, 81, 129, 101]. Conferences [175, 176, 174, 177, 178]. Conjugacy [30]. Constraint [214, 110, 110]. Constraint-Based [110]. Constructing [41]. Construction [59, 174, 53, 177, 196]. Constructions [40]. content [151]. Context [135, 212, 194]. Context-Awareness [212]. Control [180, 31, 10]. Controlled [40]. Convex [13]. Cooperative [21, 124]. Coordination [225, 225]. Cork [214]. Correcting [118, 36, 16]. Correction [17]. Cost [182]. Coventry [124]. CPAIOR [214]. Crema [111]. Crete [192]. Criterion [21]. Cryptographer [120]. Cryptography [55, 134, 155, 27, 205, 197]. Cryptology [95, 231, 88, 120, 186]. Cryptosystem [38, 39, 29]. Cryptosystems [34, 33]. CSCWD [124]. Csr [203]. CT [120]. CT-RSA [120]. Current [100]. Curve [27]. Curves [55]. Cyclic [19]. Cyclotomic [20].

Czech [75, 100].

Dagstuhl [109]. DAIS [222]. DALT [162]. DAMAS [148]. DASFAA [140]. Data [173, 82, 128, 228, 171, 83]. Database [154, 140]. Databases [184, 110]. DCOSS [223]. December [85, 89, 96, 97, 98, 95, 102, 87, 231, 88, 230, 79, 94, 186]. Decimation [43]. Decisions [143]. Declarative [93, 162]. Decoding [18]. deduction [84]. Defence [148]. Degrees [27]. Denmark [77]. Design [202, 190, 23, 124]. Designs [21, 20]. Detection [157, 119]. Determine [51]. development [138, 103]. Devices [196]. Diego [208]. Differential [28]. Digital [191]. Dimension [34]. Dimensional [32]. Directed [157]. Discovery [184, 144, 170, 173]. Discrete [25, 80]. Distributed [224, 222, 223, 79]. Diversity [21]. DNA [150, 150, 23]. DNA11 [150]. Document [130]. Documents [170]. Domain [67]. Double [56]. drawing [105]. Dresden [132]. DSV [190]. DSV-IS [190]. dual [15]. Dublin [204, 211, 212].

E-Learning [191]. e-services [90]. E4MAS [99]. EA [129]. ECCV [210]. ECIR [187]. economics [98]. EDBT [154]. Edinburgh [127, 217, 194]. Edutainment [191]. Efficient [71, 60]. Eindhoven [200]. Elliptic [27]. Embedded [1]. EMD [67]. Emerging [216]. Encryption [58, 64, 1]. ENEI [91]. Engineering [227, 176, 169, 192, 74]. Enterprise [146]. Entertainment [191]. Environments [99]. Equations [34]. Equivalence [63]. Erasure [17]. Error [118, 36, 17]. Error-Correcting [118, 36]. Errors [16]. ESOP [178]. Essays [153]. ETAPS [175, 176, 174, 177, 178]. ETRICS [216]. EURO [141]. EURO-NGI [141]. EuroGP [163]. Europe [213]. European [175, 226, 176, 110, 163, 164, 174, 187, 177, 126, 178]. EUROPKI [226]. Even [153]. Events [106]. EvoBIO [165].

EvoCOMNET [165]. EvoCOP [164]. EvoHOT [165]. EvoIASP [165]. EvoINTERACTION [165]. evolution [129]. Evolutionary [165, 164]. EvoMUSART [165]. EvoSTOC [165]. EvoWorkshops [165]. EWSN [126]. Excellence [141]. Experience [161]. Experimental [220]. Extendability [22]. Extending [56, 154, 29]. Extension [67]. Extreme [227].

Factoring [26, 63, 64]. Factoring-Based [64]. FASE [176]. FAST [125]. FATES [217]. Feature [189]. February [130, 121, 142, 118, 136, 182, 132, 120, 126, 189]. Federation [109]. Feedback [135]. FHPCN [230]. Fields [25]. Filtering [42]. Finding [48, 65, 28]. Finite [25]. Finland [227]. First [195, 98, 95, 203, 200, 87, 131, 170, 191, 156, 133, 89, 75]. FL [84]. FLOPS [193]. FMOODS [224]. FoIKS [121]. Forgery [50]. Formal [202, 202, 125, 224, 217, 132]. Forward [58]. Forward-Secure [58]. FOSSACS [175]. Foundations [94, 175, 121]. France [196, 166, 91, 142, 107, 102, 139, 129]. Francisco [223]. Frankfurt [152]. Frankfurt/Main [152]. Free [66]. Freiburg [216]. Frequency [44]. Frequency/Time [44]. Front [2, 3, 6, 11]. Frontiers [230, 103]. FSTTCS [94]. Fuji [193]. Fuji-Susono [193]. Functional [193]. Functions [26, 25, 66, 53]. Fundamental [176]. Fuzzy [111, 69].

Gabidulin [18]. Gainesville [84]. Galdana [220]. Galois [46]. Games [108]. Gan [108]. Garside [30]. GD [105]. General [48]. Generalized [21]. Generation [141]. Generator [60, 51]. Generic [59, 63]. Genetic [163]. Genus [55]. Geometric [22]. Geometry [80, 84]. Germany [110, 152, 154, 109, 123, 131, 132, 216, 228]. Gesture [139, 139]. Gibson [29]. Glasgow [135]. Goa [85]. Goal [157]. Goal-Directed [157]. Goethals [20]. Goethals-Like [20]. Good [28]. GPC [195]. GPT [29]. GRACE [230]. Graph [105]. Graz [210]. Greece [222, 192, 116]. Grid [195]. GridGIS [230]. Groups [30]. GTP [230]. Guadalajara [70]. Guess [51]. Guess-and-Determine [51]. GW [139].

Hadamard [43]. Haifa [133]. Hainan [97]. Handshakes [35]. Hangzhou [161, 191]. Harbin [104, 103]. hardware [202, 133]. Hash [67, 66, 50, 32]. HCI [210]. Held [175, 176, 174, 177, 178]. Heraklion [192]. HFE [33]. HFE-Cryptosystems [33]. HIBE [57]. Hiding [76, 62]. Hierarchy [31]. high [230, 85]. Hinterzarten [110]. HIPC [85]. HMAC [50]. hoc [87]. Homogeneous [46]. Hong [98, 101]. Hopping [44]. HPC [230]. HPC-GTP [230]. HSCC [180]. Human [139, 210, 200]. Human-Computer [139, 210]. Hungary [163, 121, 164, 165]. Hybrid [180]. Hyderabad [113, 114, 94].

ICA [147]. ICB [101]. ICCS [215]. ICDM [228]. ICFCA [132]. ICISC [186]. ICSE [104]. ICSOC [96]. ID [37]. ID-Based [37]. Identity [59]. Identity-Based [59]. IEEE [223, 208]. IFIP [209, 181, 222, 224, 116, 201]. IFIP-TC [209]. IH [76]. II [124, 99, 114]. IICS [70, 166]. III [162]. image [228]. Imai [34]. Immune [183]. Immunity [53]. Implementation [107]. Improved [46, 49]. Independent [147]. India [85, 88, 113, 114, 79, 94]. Indifferentiable [66]. INDOCRYPT [88]. Inductive [184, 110]. Industrial [228]. Informatics [172, 145, 208]. Information [115, 231, 187, 216, 78, 121, 76, 161, 95, 86, 186]. Infrastructure [226]. Innovative [70, 166]. Inspired [115]. Integer [26]. Integration [192, 214]. Intelligence [208, 143, 172, 208].

Intelligent [136]. Interaction [139, 210, 127]. Interactive [190]. Interlace [45]. Interleaved [17]. International [227, 175, 215, 183, 167, 162, 176, 76, 196, 214, 111, 209, 70, 184, 122, 206, 144, 106, 166, 91, 198, 141, 172, 161, 195, 225, 202, 185, 98, 135, 125, 181, 77, 222, 137, 92, 204, 118, 112, 211, 136, 69, 223, 139, 190, 224, 152, 217, 203, 192, 193, 212, 93, 108, 174, 180, 159, 200, 115, 154, 109, 87, 160, 231, 140, 182, 131, 88, 208, 132, 216, 177, 117, 170, 82, 79, 191, 155, 94, 147, 194, 124, 149, 148, 158, 143, 156, 133, 201, 179, 199, 207, 99, 186, 168, 197, 220]. international [151, 85, 89, 96, 127, 78, 130, 73, 90, 150, 97, 121, 74, 107, 102, 105, 84, 123, 81, 75, 230, 104, 86, 116, 129, 119, 101, 205]. Internet [70, 206, 166, 141, 98]. Interoperable [196, 222]. Interpolation [26]. Interpolations [27]. Interpretation [117, 45]. Introduction [7]. intrusion [119]. Invertible [47]. Invited [184, 122, 112, 149, 148, 158, 99]. IR [187]. Ireland [214, 204, 211, 212, 105]. Irreducible [19, 4]. ISAAC [97]. ISBMDA [82]. ISDF [230]. ISI [208]. Island [139, 86]. ISNN [207]. ISPA [230]. ISPEC [161]. Israel [108, 133]. Issues [169]. ISVC [89]. Italian [183, 218]. Italy [183, 167, 226, 151, 111, 218, 141, 225, 202, 69, 224, 160, 230]. Itatiaia [199]. IV [169, 188]. IWDC [79]. IWSN [104].

Jamaica [106]. January [93, 115, 117, 113, 114, 104, 100, 101, 103]. Japan [80, 193, 115]. Japanese [80]. JCDCG [80]. Jeju [86]. Joint [175, 176, 174, 177, 178]. Josas [102]. Jose [120]. Journal [128]. Jouy [102]. Jouy-en-Josas [102]. July [162, 213, 127, 122, 78, 141, 135, 125, 112, 190, 217, 108, 123, 228, 68, 194, 149, 148, 156, 99]. June [227, 183, 226, 76, 214, 213, 70, 166, 150, 225, 77, 222, 107, 223, 224, 203, 81, 216, 158, 207].

KDID [184]. KDLL [144]. KDXD [170]. Key [226, 50, 38, 31, 197]. Key-Recovery [50]. Keys [58]. Kharagpur [79]. Kingston [155]. Kleene [182]. Knowledge [173, 121, 123, 184, 144, 170]. Kong [98, 101]. Korea [86, 186].

Lake [89]. LAMAS [156]. Landscape [131]. Language [199]. Languages [162, 225, 93, 178, 68]. Large [44, 81]. Large-scale [81]. Latent [189]. Latin [145, 145]. LCPC [68]. Learning [191, 127, 156]. lectures [202]. Leipzig [131, 228]. Level [43]. Life [144]. Like [20, 18]. Lille [129]. Limerick [105]. Linear [92, 42, 44, 22, 5]. linearity [25]. Linearization [34]. Linguistics [136]. Literature [144]. LOCA [212]. Locally [47]. Location [212]. Location- [212]. Logarithm [25]. Logic [111, 69, 193, 158, 159]. Logical [213]. London [150, 159, 187, 158]. LOPSTR [159]. lSSC [81]. Lyngby [77].

MABS [149]. Machine [127]. Magyarik [39]. Main [152]. Mallorca [137]. Management [91, 123]. Map [131]. Maps [131, 37]. March [175, 176, 196, 110, 145, 152, 134, 174, 180, 154, 177, 147, 178, 179, 205]. Mare [183]. marketing [228]. Marseille [142]. MARTES [106]. Mathematical [123]. Mathematics [168]. Matrices [47, 4]. Matrix [12]. Matsumoto [34]. Matter [14, 2, 3, 6, 11]. Maximum [53]. May [215, 220, 214, 209, 206, 198, 218, 195, 202, 74, 204, 211, 139, 212, 210, 200, 109, 75, 208, 124, 201, 199, 207]. MD [68]. MDAI [143]. MDD [106]. Medical [82]. medicine [228]. MEGA [104]. Membrane [112]. Memory [153]. Menorca [220]. Merin [100]. Message [62]. Messages [37]. Method [55, 41]. Methodology [83]. Methods [224, 182, 202]. Mexico [70, 136].

Micropayment [32]. Mining [110, 171, 173, 83, 228]. MKM [123]. MLMI [127]. Mobile [209, 87]. MODAUI [106]. Model [117, 8, 62, 179]. Modeling [160, 143, 194]. MoDELS [106, 198, 225]. MoDeVA [106]. Modular [52]. Molecular [167]. Montego [106]. motion [75]. MRC [194]. MSN [87]. MTiP [106]. Multi [67, 122, 169, 32, 148, 158, 156, 37, 99, 149]. Multi-Agent [122, 169, 148, 158, 156, 99]. Multi-Agent-Based [149]. Multi-Dimensional [32]. Multi-Messages [37]. Multi-Property-Preserving [67]. Multimedia [135]. multimodal [127]. Multiplication [56]. Multisignature [37]. Multivariate [65, 47]. Munich [154]. Music [160].

NAIS [183]. Nancy [91]. Naples [69]. Natural [183]. Nelson [130]. Netherlands [162, 96, 122, 78, 200, 149, 148, 156, 99]. Nets [183]. Network [141, 8, 98, 104]. Networking [209, 230, 209]. Networks [209, 87, 9, 219, 207, 126]. Neural [183, 207, 183]. Newcastle [125, 190]. Next [141]. NfC [106]. NGI [141]. Nice [196]. NMAC [50]. NOLISP [92]. Non [25, 92]. Non-Linear [92]. Non-linearity [25]. Nonlinear [42, 92]. Nonnegative [5]. Norway [90, 146, 205]. Novel [41]. November [82, 133]. NV [89, 118]. NY [134, 201, 197].

OAEP [61]. Object [157, 224]. Object-Based [224]. OCLWS [106]. October [80, 184, 106, 137, 69, 131, 116, 129]. OFDM [40]. One [64, 45]. One-Wayness [64]. Online [137, 71]. Open [24, 224]. Optimization [214, 164, 189, 71]. Oracle [57]. oriented [138, 96]. Osaka [115]. Oulu [227].

PA [73]. Pacific [229, 173, 103]. Padding [66]. PADL [93]. PAKDD [144, 171, 173]. Palma [137]. Papers [205, 80, 183, 162, 76, 196, 111, 70, 184, 122, 110, 106, 166, 91, 135, 125, 77, 146, 137, 92, 112, 69, 139, 190, 217, 192, 108, 159, 109, 160, 182, 131, 155, 194, 189, 124, 149, 148, 158, 156, 133, 99, 186, 168, 151, 127, 78, 73, 90, 150, 141, 74, 107, 102, 105, 84, 123, 81, 75, 68, 86, 116, 129, 119]. PARA [77]. Parallel [77, 37, 68, 168]. Parameters [57]. ParDMCom [230]. Paris [166]. Park [68]. Part [175, 215, 176, 174, 177, 113, 114, 178, 207]. Partial [50]. PATAT [73]. Patterns [28]. PDCE [230]. Perfect [41]. Performance [209, 85, 230]. Perron [4]. Perspectives [189]. Persuasive [200, 200]. PERVASIVE [204, 195, 185, 211, 204, 211]. Petersburg [203]. Pisa [160]. Pittsburgh [73]. PKI [226]. Plaintext [62]. Poland [168]. Polynomial [47, 27]. Polynomials [38, 65, 20, 45]. Polyphase [41]. Popular [66]. Porto [184]. Portugal [209, 184, 82]. Portuguese [199]. Positive [5]. Power [40, 10]. Power-Controlled [40]. Poznan [168]. PPAM [168]. PPK [197]. Practical [169, 93]. Practice [175, 226, 176, 161, 174, 177, 178, 100, 197, 73]. Prague [75]. Prefix [66]. Prefix-Free [66]. Preserving [67]. Private [58]. Problem [26, 38, 30, 9]. Problems [214, 24, 71]. Proceedings [215, 113, 114, 143, 207, 227, 175, 167, 226, 176, 214, 209, 206, 144, 198, 218, 172, 161, 195, 225, 185, 163, 145, 181, 142, 222, 95, 204, 118, 211, 136, 223, 224, 164, 152, 203, 193, 134, 212, 93, 174, 180, 210, 200, 87, 231, 187, 140, 171, 88, 208, 132, 216, 177, 117, 170, 173, 82, 79, 191, 94, 147, 165, 178, 201, 179, 199, 100, 197, 220, 85, 89, 213, 96, 130, 97, 98, 121, 229, 230, 228, 120, 126, 104, 101, 103]. Process [91]. Processes [227]. Processing [92, 136, 199, 168, 151]. Program [159]. Programming [214, 163, 193, 178, 227, 122, 178]. programs [102]. progress [71, 88]. ProMAS [122]. proofs [102]. Properties [59]. Property [67]. PROPOR [199]. Protocols [209]. Provable [60]. Provers [72]. Pseudorandom [60]. Public [57, 38, 197, 226, 197]. Public-Key [38, 197].

RAID [119]. Ramat [108]. Ramat-Gan [108]. Random [57]. Rank [21, 16]. Rapid [192]. Reading [215]. Realizations [43]. RECOMB [167]. Reconstructing [38]. Recovery [50]. Reduction [30]. Reed [17]. Regarding [63]. Register [42]. Related [26, 25, 24, 27]. Relation [36]. Relational [182, 182]. Relationship [62]. representation [151]. Republic [75, 100]. Research [181, 169, 187, 131, 221, 74, 103, 167]. Resource [9, 219]. Results [48, 13, 12]. Retrieval [135, 160, 187, 194]. Revised [80, 183, 162, 76, 196, 111, 70, 184, 122, 110, 106, 166, 91, 135, 125, 77, 146, 137, 92, 112, 69, 139, 190, 217, 192, 108, 159, 109, 160, 182, 131, 155, 194, 189, 124, 149, 148, 158, 156, 133, 99, 186, 168, 205, 151, 127, 78, 73, 90, 150, 141, 74, 107, 102, 105, 84, 123, 81, 75, 68, 86, 116, 129, 119]. Ring [63]. Rings [46]. RISE [192]. Rome [218]. Root [4]. Roots [65]. RSA [120, 65, 63, 60, 35]. RSA-Based [60, 35]. Russia [203].

S [230]. S-GRACE [230]. SAC [155]. Safe [196]. San [223, 208, 120]. Santa [180]. Sanya [97]. Sardinia [151]. Satellite [106]. Scalable [1]. Scalar [56]. scale [81]. Scheme [31]. Schemes [32, 36, 37]. School [202]. Science [175, 215, 144, 142, 153, 203, 182, 94, 100]. Scientific [77, 81]. SCVMA [75]. SEA [1]. Search [49, 157]. Searchable [58]. Seattle [119]. Second [196, 213, 141, 161, 223, 192, 212, 115, 194, 15, 116, 99, 127, 74]. Secret [36, 35]. Secure [58, 196, 54]. Security [61, 66, 172, 161, 125, 95, 231, 208, 216, 64, 60, 186, 52, 86, 185]. Selected [80, 183, 162, 76, 196, 111, 184, 110, 106, 91, 135, 125, 77, 146, 137, 92, 112, 69, 139, 217, 192, 159, 109, 182, 131, 155, 194, 189, 124, 158, 156, 133, 99, 186, 168, 205, 151, 127, 78, 73, 90, 150, 141, 74, 107, 102, 123, 86, 116, 129, 155]. Selection [189]. Self [15, 51]. Self-dual [15]. Self-Shrinking [51]. Semantic [221]. Semantics [128]. Seminar [182]. Sensor [223, 87, 126]. Seoul [186]. Separation [147]. September [151, 111, 91, 90, 192, 105, 159, 84, 229, 160, 119, 168]. Sequences [42, 44, 40, 43, 41]. SERA [74]. Series [37]. Series-Parallel [37]. Service [96]. Service-oriented [96]. Services [209, 90]. Seventeen [72]. SFM [202]. SHA [48, 28, 49]. SHA-0 [49]. SHA-1 [48, 28]. Shanghai [229, 231]. Sharing [36]. Shift [42]. Shift-Register [42]. Shimon [153]. Short [58, 57]. Shrinking [51]. Signal [147, 228]. Signatures [59]. Simulation [139, 149]. Singapore [144, 172, 140, 171, 170, 173]. Sketch [54]. SKLOIS [95]. Slovenia [189]. SLSPS [189]. Small [1]. Smart [196, 181]. SOFSEM [100]. Software [227, 175, 176, 202, 217, 192, 174, 177, 94, 178, 133, 179, 138, 74, 169]. Solomon [17]. Solution [5]. Some [13, 12]. Sophia [107]. Sorrento [230]. Sozopol [81]. Space [21]. Space-Time [21]. Spain [220, 76, 181, 137, 92, 143]. Sparsity [25]. Spatial [75]. SPC [185]. Special [43]. Specification [190]. Spectral [45]. Speech [92]. SPIN [179]. St [203, 182]. STACS [142]. Standard [62]. State [77]. Statistical [189]. Strategy [65]. Stream [24, 52]. Structure [189]. Structures [175, 30]. Subspace [189]. sul [183]. Sums [46]. Support [15]. Supported [124]. Susono [193]. Swansea [213]. Switzerland [206, 126]. Symmetric [16]. Symposium [145, 142, 118, 203, 193, 93, 159, 160, 82, 178, 207, 89, 97, 121, 105, 119]. Synthesis [159]. System [157, 5]. Systems [183, 209, 70, 122, 166, 141, 202, 222, 33, 169, 223, 190, 224, 152, 174, 180, 140, 188, 178, 148, 158, 156, 201, 99, 78, 130, 121, 229].

TACAS [174]. Tahoe [89]. Taichung [195]. Taiwan [195]. TAMC [198]. Tarragona [181, 143]. Tarski [182]. TC [209, 201]. TCC [134]. TEAA [146]. Techniques [214, 192, 83]. Technologies [162, 209, 104, 90, 191]. Technology [200, 115, 154, 94]. Templates [54]. Ternary [22]. TES [90]. TESTCOM [201]. Testing [217, 133, 201]. Text [136]. Theoretical [145, 142, 153, 94]. Theory [175, 226, 176, 203, 134, 174, 231, 177, 178, 219, 100, 83, 197, 73, 198, 134]. Third [226, 162, 214, 122, 198, 185, 135, 125, 137, 134, 160, 143, 207, 96, 126]. Time [21, 44]. timetabling [73]. TMRA [131]. Tokyo [80]. Tools [174]. Topic [131]. Topics [120]. Track [120]. Trading [64]. Transactions [138, 188]. Transform [67, 43]. Transformation [159]. Transformations [33]. Transmission [40]. Tree [31]. Trends [216, 100, 146]. Trondheim [90, 146]. Trust [125]. Turin [226]. Two [25, 45, 30, 43]. Two-Level [43]. Two-Variable [45]. Tyne [125, 190]. TYPES [102, 102].

UK [215, 213, 127, 185, 135, 125, 190, 217, 159, 187, 194, 124, 158]. upon [125, 190]. UPWN [230]. USA [89, 73, 74, 118, 223, 134, 93, 180, 84, 208, 117, 120, 68, 147, 201, 119, 197]. User [135]. Using [56, 50, 30]. Utrecht [162, 122, 149, 148, 156, 99].

V [73, 128]. Valdivia [145]. Variable [45]. Variables [53]. Variants [65, 20]. Vegas [118]. Venice [167]. Verifiable [36]. Verification [190, 133, 202, 117]. VI [149, 158]. Vienna [175, 176, 112, 174, 177, 178, 179]. Vietri [183]. Vigoni [141]. VII [130]. Villa [141].

Vision [210, 113, 114]. VISUAL [78, 157, 89, 75, 151, 78]. VLBV [151]. VLDB [146]. VMCAI [117]. VOCUS [157].

WA [119]. WAC [116]. Wagner [39]. WAOA [137]. Wayness [64]. WCC [205]. WEA [220]. Web [109, 104, 103, 228, 221]. Weights [19, 46, 15]. Weil [46]. Welch [18]. Well [200]. Well-Being [200]. WG [181, 222, 224, 201]. WILF [111, 69]. Wine [98]. Wired [206]. Wired/Wireless [206]. Wireless [209, 219, 206, 141, 126]. WIRN [183]. WISA [86]. WISI [172]. WiSME [106]. Without [57]. WMC [112]. WOMP [230]. Work [124]. Workshop [183, 226, 162, 76, 196, 111, 70, 184, 122, 110, 144, 166, 141, 172, 135, 125, 77, 146, 137, 112, 69, 139, 190, 217, 192, 212, 180, 210, 115, 109, 171, 182, 131, 170, 79, 155, 194, 189, 149, 148, 158, 156, 179, 199, 99, 205, 220, 151, 127, 130, 90, 150, 98, 102, 84, 75, 68, 126, 86, 116]. Workshops [106, 91, 230, 104]. World [72]. WSCOBPM [91]. Wuhan [87]. WUsCAM [106]. WWIC [206]. WWW [103].

XHPC [230]. XML [170]. XP [227]. XRA [104].

York [185, 134, 201, 197].

Z4 [20]. Zealand [130]. Zürich [126].

References

Standaert:2006:SSE

[1] François-Xavier Standaert, Gilles Piret, Neil Gershenfeld, and Jean-Jacques Quisquater. SEA: A scalable encryption algorithm for small embedded applications. Lecture Notes in Computer Science, 3928:222–236, 2006. ISBN 3-540-33311-8 (print), 3-540-33312-6 (e-

book). URL http://link.springer.
com/chapter/10.1007/11733447_16.

Stanczak:2006:FMa

[2] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Front matter. Lecture Notes in Computer Science, 4000:??, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-
3-540-46249-1/1.

Stanczak:2006:FMb

[3] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Front matter. Lecture Notes in Computer Science, 4000:1, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-
3-540-46249-1/1/1.

Stanczak:2006:CPR

[4] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Chapter 1: On the Perron root of irreducible matrices. Lecture Notes in Computer Science, 4000:3–49, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/chapter/10.1007/11818762_1/.

Stanczak:2006:CPS

[5] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Chapter 2: On the positive solution to a linear system with nonnegative coefficients. Lecture Notes in Computer Science, 4000: 51–68, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.

                                        com/accesspage/chapter/10.1007/
                                        11818762_1;   http://link.springer.
                                        com/chapter/10.1007/11818762_2/.

Stanczak:2006:FMc

[6] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Front matter. Lecture Notes in Computer Science, 4000:69, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/chapter/10.1007/
                                        11818762_2;   http://link.springer.
                                        com/content/pdf/bfm:978-3-540-46249-
                                        1/2/1.

Stanczak:2006:CI

[7] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Chapter 3: Introduction. Lecture Notes in Computer Science, 4000:71–73, 2006. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/chapter/
                                        10.1007/11818762_3/.

Stanczak:2006:CNM

[8] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Chapter 4: Network model. Lecture Notes in Computer Science, 4000:75–89, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/chapter/10.1007/
                                        11818762_3;   http://link.springer.
                                        com/chapter/10.1007/11818762_4/.

Stanczak:2006:CRA

[9] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Chapter 5: Resource allocation problem in communications networks. Lecture Notes in Computer Science, 4000:91–128, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/accesspage/chapter/10.1007/
11818762_4;   http://link.springer.
com/chapter/10.1007/11818762_5/.

Stanczak:2006:CPC

[10] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Chapter 6: Power control algorithm. Lecture Notes in Computer Science, 4000:129–152, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/accesspage/chapter/10.1007/
11818762_5;   http://link.springer.
com/chapter/10.1007/11818762_6/.

Stanczak:2006:FMd

[11] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Front matter. Lecture Notes in Computer Science, 4000:153, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/accesspage/chapter/10.1007/
11818762_6;   http://link.springer.
com/content/pdf/bfm:978-3-540-46249-
1/3/1.

Stanczak:2006:ASC

[12] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Appendix A: Some concepts and results from matrix analysis. Lecture Notes in Computer Science, 4000:155–170, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/chapter/10.1007/11818762_7/.

Stanczak:2006:ABS

[13] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Appendix B: Some concepts and results from convex analysis. Lecture Notes in Computer Science, 4000:171–183, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/chapter/10.1007/
                                        11818762_7;   http://link.springer.
                                        com/chapter/10.1007/11818762_8/.

Stanczak:2006:BM

[14] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche. Back matter. Lecture Notes in Computer Science, 4000: ??, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/chapter/10.1007/
                                        11818762_8;   http://link.springer.
                                        com/content/pdf/bbm:978-3-540-46249-
                                        1/1.

Shiromoto:2006:SSW

[15] Keisuke Shiromoto. Second support weights for binary self-dual codes. In Ytrehus [205], pages 1–13. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Pilipchuk:2006:CCS

[16] Nina I. Pilipchuk and Ernst M. Gabidulin. On codes correcting symmetric rank errors. In Ytrehus [205], pages 14–21. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Schmidt:2006:EEC

[17] Georg Schmidt, Vladimir R. Sidorenko, and Martin Bossert. Error and erasure correction of interleaved Reed– Solomon codes. In Ytrehus [205], pages 22–35. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Loidreau:2006:WLA

[18] Pierre Loidreau. A Welch–Berlekamp like algorithm for decoding Gabidulin codes. In Ytrehus [205], pages 36– 45. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Aubry:2006:WBI

[19] Yves Aubry and Philippe Langevin. On the weights of binary irreducible cyclic codes. In Ytrehus [205], pages 46–54. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Lahtonen:2006:DZG

[20] Jyrki Lahtonen, Kalle Ranto, and Roope Vehkalahti. 3-designs from Z4–Goethalslike codes and variants of cyclotomic polynomials. In Ytrehus [205], pages 55–68. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Hammons:2006:STC

[21] A. Roger Hammons, Jr. Space-time code designs based on the generalized binary rank criterion with applications to cooperative diversity. In Ytrehus [205], pages 69–84. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Maruta:2006:GCE

[22] Tatsuya Maruta and Kei Okamoto. Geometric conditions for the extendability of ternary linear codes. In Ytrehus [205], pages 85–99. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Milenkovic:2006:DCD

[23] Olgica Milenkovic and Navin Kashyap. On the design of codes for DNA computing. In Ytrehus [205], pages 100– 119. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Canteaut:2006:OPR

[24] Anne Canteaut. Open problems related to algebraic attacks on stream ciphers. In Ytrehus [205], pages 120–134. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Brandstatter:2006:NLS

[25] Nina Brandstätter, Tanja Lange, and Arne Winterhof. On the non-linearity and sparsity of Boolean functions related to the discrete logarithm in finite fields of characteristic two. In Ytrehus [205], pages 135–143. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Adelmann:2006:IFR

[26] Clemens Adelmann and Arne Winterhof. Interpolation of functions related to the integer factoring problem. In Ytrehus [205], pages 144–154. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Satoh:2006:DPI

[27] Takakazu Satoh. On degrees of polynomial interpolations related to elliptic curve cryptography. In Ytrehus [205], pages 155–163. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Matusiewicz:2006:FGD

[28] Krystian Matusiewicz and Josef Pieprzyk. Finding good differential patterns for attacks on SHA-1. In Ytrehus [205], pages 164–177. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Overbeck:2006:EGA

[29] Raphael Overbeck. Extending Gibson’s attacks on the GPT cryptosystem. In Ytrehus [205], pages 178–188. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Samuel:2006:RCP

[30] Maffre Samuel. Reduction of conjugacy problem in braid groups, using two Garside structures. In Ytrehus [205], pages 189–201. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Santis:2006:NKA

[31] Alfredo De Santis, Anna Lisa Ferrara, and Barbara Masucci. A new key assignment scheme for access control in a complete tree hierarchy. In Ytrehus [205], pages 202–217. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Nguyen:2006:MDH

[32] Quan Son Nguyen. Multi-dimensional hash chains and application to micropayment schemes. In Ytrehus [205], pages 218–228. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Felke:2006:ATH

[33] Patrick Felke. On the affine transformations of HFE-cryptosystems and systems with branches. In Ytrehus [205], pages 229–241. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Diene:2006:DLE

[34] Adama Diene, Jintai Ding, Jason E. Gower, Timothy J. Hodges, and Zhijun Yin. Dimension of the linearization equations of the Matsumoto–Imai cryptosystems. In Ytrehus [205], pages 242–251. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Vergnaud:2006:RBS

[35] Damien Vergnaud. RSA-based secret handshakes. In Ytrehus [205], pages 252–274. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Nikov:2006:RBV

[36] Ventzislav Nikov and Svetla Nikova. On a relation between verifiable secret sharing schemes and a class of errorcorrecting codes. In Ytrehus [205], pages 275–290. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Wang:2006:IBS

[37] Lihua Wang, Eiji Okamoto, Ying Miao, Takeshi Okamoto, and Hiroshi Doi. ID-based series-parallel multisignature schemes for multi-messages from bilinear maps. In Ytrehus [205], pages 291–303. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Faure:2006:NPK

[38] Cédric Faure and Pierre Loidreau. A new public-key cryptosystem based on the problem of reconstructing ppolynomials. In Ytrehus [205], pages 304–315. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Levy-dit-Vehel:2006:WC

[39] Françoise Levy dit Vehel and Ludovic Perret. On the Wagner–Magyarik cryptosystem. In Ytrehus [205], pages 316– 329. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Schmidt:2006:CCS

[40] Kai-Uwe Schmidt and Adolf Finger. Constructions of complementary sequences for power-controlled OFDM transmission. In Ytrehus [205], pages 330–345. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Zeng:2006:NMC

[41] Xiangyong Zeng, Lei Hu, and Qingchong Liu. A novel method for constructing almost perfect polyphase sequences. In Ytrehus [205], pages 346–353. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Gammel:2006:LFN

[42] Berndt M. Gammel and Rainer Göttfert. Linear filtering of nonlinear shift-register sequences. In Ytrehus [205], pages 354–370. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Yu:2006:RDH

[43] Nam Yul Yu and Guang Gong. Realizations from decimation Hadamard transform for special classes of binary sequences with two-level autocorrelation. In Ytrehus [205], pages 371–385. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Hong:2006:FTH

[44] Yun-Pyo Hong and Hong-Yeop Song. Frequency/time hopping sequences with large linear complexities. In Ytrehus [205], pages 386–396. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Riera:2006:OTV

[45] Constanza Riera and Matthew G. Parker. One and two-variable interlace polynomials: a spectral interpretation. In Ytrehus [205], pages 397–411. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Ling:2006:IBW

[46] San Ling and Ferruh Özbudak. Improved bounds on Weil sums over Galois rings and homogeneous weights. In Ytrehus [205], pages 412–426. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Lobo:2006:LIM

[47] Ruben G. Lobo, Donald L. Bitzer, and Mladen A. Vouk. Locally invertible multivariate polynomial matrices. In Ytrehus [205], pages 427–442. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

DeCanniere:2006:FSC

[48] Christophe De Cannière and Christian Rechberger. Finding SHA-1 characteristics: General results and applications. Lecture Notes in Computer Science, 4284:1–20, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/11935230_1.pdf;   http:/
/link.springer.com/content/pdf/
bfm:978-3-540-49476-8/1.pdf.

Naito:2006:ICS

[49] Yusuke Naito, Yu Sasaki, Takeshi Shimoyama, Jun Yajima, Noboru Kunihiro, and Kazuo Ohta. Improved collision search for SHA-0. Lecture Notes in Computer Science, 4284:21–36, 2006. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/11935230_2.pdf.

Contini:2006:FPK

[50] Scott Contini and Yiqun Lisa Yin. Forgery and partial key-recovery attacks on HMAC and NMAC using hash collisions. Lecture Notes in Computer Science, 4284:37–53, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
11935230_3.pdf.

Zhang:2006:NGD

[51] Bin Zhang and Dengguo Feng. New guess-and-determine attack on the selfshrinking generator. Lecture Notes in Computer Science, 4284:54–68, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/11935230_4.pdf.

Paul:2006:SSC

[52] Souradyuti Paul and Bart Preneel. On the (in)security of stream ciphers based on arrays and modular addition. Lecture Notes in Computer Science, 4284:69–83, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/11935230_
5.pdf.

Li:2006:CAB

[53] Na Li and Wen-Feng Qi. Construction and analysis of Boolean functions of 2t + 1 variables with maximum algebraic immunity. Lecture Notes in Computer Science, 4284:84–98, 2006. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/11935230_6.pdf.

Li:2006:SSB

[54] Qiming Li, Yagiz Sutcu, and Nasir Memon. Secure sketch for biometric templates. Lecture Notes in Computer Science, 4284:99–113, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        11935230_7.pdf.

Gaudry:2006:ACM

[55] P. Gaudry, T. Houtmann, D. Kohel, C. Ritzenthaler, and A. Weng. The 2-adic CM method for genus 2 curves with application to cryptography. Lecture Notes in Computer Science, 4284: 114–129, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/11935230_
                                        8.pdf.

Avanzi:2006:ESM

[56] Roberto Avanzi, Vassil Dimitrov, Christophe Doche, and Francesco Sica. Extending scalar multiplication using double bases. Lecture Notes in Computer Science, 4284:130–144, 2006. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/11935230_9.pdf.

Chatterjee:2006:HSP

[57] Sanjit Chatterjee and Palash Sarkar. HIBE with short public parameters without random oracle. Lecture Notes in Computer Science, 4284:145–160, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/11935230_10.pdf.

Attrapadung:2006:FSS

[58] Nuttapong Attrapadung, Jun Furukawa, and Hideki Imai. Forward-secure and searchable broadcast encryption with short ciphertexts and private keys. Lecture Notes in Computer Science, 4284: 161–177, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/11935230_
11.pdf.

Galindo:2006:GCI

[59] David Galindo, Javier Herranz, and Eike Kiltz. On the generic construction of identity-based signatures with additional properties. Lecture Notes in Computer Science, 4284:178–193, 2006. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/11935230_12.pdf.

Steinfeld:2006:PSE

[60] Ron Steinfeld, Josef Pieprzyk, and Huaxiong Wang. On the provable security of an efficient RSA-based pseudorandom generator. Lecture Notes in Computer Science, 4284:194–209, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/content/
pdf/10.1007/11935230_13.pdf.

Boldyreva:2006:SO

[61] Alexandra Boldyreva and Marc Fischlin. On the security of OAEP. Lecture Notes in Computer Science, 4284: 210–225, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/11935230_
14.pdf.

Teranishi:2006:RBS

[62] Isamu Teranishi and Wakaha Ogata. Relationship between standard model plaintext awareness and message hiding. Lecture Notes in Computer Science, 4284:226–240, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/11935230_
                                        15.pdf.

Leander:2006:ERF

[63] Gregor Leander and Andy Rupp. On the equivalence of RSA and factoring regarding generic ring algorithms. Lecture Notes in Computer Science, 4284: 241–251, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/11935230_
                                        16.pdf.

Paillier:2006:TOW

[64] Pascal Paillier and Jorge L. Villar. Trading one-wayness against chosenciphertext security in factoring-based encryption. Lecture Notes in Computer Science, 4284:252–266, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        11935230_17.pdf.

Jochemsz:2006:SFR

[65] Ellen Jochemsz and Alexander May. A strategy for finding roots of multivariate polynomials with new applications in attacking RSA variants. Lecture Notes in Computer Science, 4284:267–282, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.com/content/
                                        pdf/10.1007/11935230_18.pdf.

Chang:2006:ISA

[66] Donghoon Chang, Sangjin Lee, Mridul Nandi, and Moti Yung. Indifferentiable security analysis of popular hash functions with prefix-free padding. Lecture Notes in Computer Science, 4284: 283–298, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/11935230_
19.pdf.

Bellare:2006:MPP

[67] Mihir Bellare and Thomas Ristenpart. Multi-property-preserving hash domain extension and the EMD transform. Lecture Notes in Computer Science, 4284: 299–314, 2006. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/11935230_
20.pdf.

Pugh:2006:LCP

[68] William Pugh and Chau-Wen Tseng, editors. Languages and compilers for parallel computing: 15th workshop, LCPC 2002, College Park, MD, USA, July 25– 27, 2002, revised papers, volume 2481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30781-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 2002. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=2481.

Gesu:2006:FLA

[69] Vito Gesù, Francesco Masulli, and Alfredo Petrosino, editors. Fuzzy Logic and Applications: 5th International Workshop, WILF 2003, Naples, Italy, October 9–11, 2003. Revised Selected Papers, volume 2955 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31019-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=2955.

Boehme:2006:IIC

[70] Thomas Boehme, Victor M. Larios Rosillo, and Herwig Unger, editors. Innovative Internet Community Systems: 4th International Workshop, IICS 2004, Guadalajara, Mexico, June 21– 23, 2004. Revised Papers, volume 3473 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-28880-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3473.

Bampis:2006:EAO

[71] Evripidis Bampis, Klaus Jansen, and Claire Kenyon. Efficient approximation and online algorithms: recent progress on classical combinatorical optimization problems and new applications, volume 3484 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32213-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3484.

Wiedijk:2006:SPW

[72] Freek Wiedijk, editor. The Seventeen Provers of the World, volume 3600 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-30704-4. ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 157 pp. LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3600. Foreword by

Dana S. Scott.

Burke:2006:PTA

[73] Edmund Burke and Michael Trick, editors. Practice and theory of automated timetabling V: 5th international conference, PATAT 2004, Pittsburgh, PA, USA, August 18–20, 2004, revised selected papers, volume 3616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30705-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3616.

Dosch:2006:SER

[74] Walter Dosch, Roger Y. Lee, and Chisu Wu, editors. Software engineering research and applications: second international conference, SERA 2004, Los Angeles, CA, USA, May 5–7, 2004, revised selected papers, volume 3647 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32133-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.

                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3647.

MacLean:2006:SCV

[75] W. James MacLean, editor. Spatial coherence for visual motion analysis: first international workshop, SCVMA 2004, Prague, Czech Republic, May 15, 2004. revised papers, volume 3667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32533-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3667.

Barni:2006:IHI

[76] Mauro Barni, Jordi Herrera-Joancomartí, Stefan Katzenbeisser, and Fernando Pérez-González, editors. Information Hiding: 7th International Workshop, IH 2005, Barcelona, Spain, June 6–8, 2005. Revised Selected Papers, volume 3727 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-29039-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3727.

Dongarra:2006:APC

[77] Jack Dongarra, Kaj Madsen, and Jerzy Wasniewski, editors. Applied Parallel Computing: State of the Art in Scientific Computing. 7th International Workshop, PARA 2004, Lyngby, Denmark, June 20–23, 2004. Revised Selected Papers, volume 3732 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-29067-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3732.

Bres:2006:VII

[78] Stéphane Bres and Robert Laurini, editors. Visual information and information systems: 8th international conference, VISUAL 2005, Amsterdam, The Netherlands, July 5, 2005: revised selected papers, volume 3736 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30488-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.I6 V558 2005. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3736.

Pal:2006:DCI

[79] Ajit Pal, Ajay D. Kshemkalyani, Rajeev Kumar, and Arobinda Gupta, editors. Distributed Computing — IWDC 2005: 7th International Workshop, Kharagpur, India, December 27–30, 2005. Proceedings, volume 3741 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30959-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3741.

Akiyama:2006:DCG

[80] Jin Akiyama, Mikio Kano, and Xuehou Tan, editors. Discrete and Computational Geometry: Japanese Conference, JCDCG 2004, Tokyo, Japan, October 8–11, 2004, Revised Selected Papers, volume 3742 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30467-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3742.

Lirkov:2006:LSS

[81] Ivan Lirkov, Svetozar Margenov, and Jerzy Waśniewski, editors. Largescale scientific computing: 5th international conference, lSSC 2005, Sozopol, Bulgaria, June 6–10, 2005, revised papers, volume 3743 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31994-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3743.

Oliveira:2006:BMD

[82] José Luís Oliveira, Víctor Maojo, Fernando Martin-Sanchez, and António Sousa Pereira, editors. Biological and Medical Data Analysis: 6th International Symposium, ISBMDA 2005, Aveiro, Portugal, November 10–11, 2005. Proceedings, volume 3745 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-29674-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3745.

Williams:2006:DMT

[83] Graham J. Williams and Simeon J. Simoff, editors. Data Mining: Theory, Methodology, Techniques, and Applications, volume 3755 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32547-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3755.

Hong:2006:ADG

[84] Hoon Hong and Dongming Wang, editors. Automated deduction in geometry: 5th international workshop, ADG 2004, Gainesville, FL, USA, September 16– 18, 2004: revised papers, volume 3763 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31332-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA448.D38 I577 2004. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3763.

Bader:2006:HPC

[85] David A. Bader, Manish Parashar, Varadarajan Sridhar, and Viktor K. Prasanna, editors. High performance computing — HIPC 2005: 12th international conference, Goa, India, December 18–21, 2005. proceedings, volume 3769 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30936-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3769.

Song:2006:ISA

[86] JooSeok Song, Taekyoung Kwon, and Moti Yung, editors. Information security applications: 6th international workshop, WISA 2005, Jeju Island, Korea, August 22–24, 2005: revised selected papers, volume 3786 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31012-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25 W577 2006. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-9743&
                                        volume=3786.

Jia:2006:MAH

[87] Xiaohua Jia, Jie Wu, and Yanxiang He, editors. Mobile Ad-hoc and Sensor Networks: First International Conference, MSN 2005, Wuhan, China, December 13–15, 2005. Proceedings, volume 3794 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30856-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3794.

Maitra:2006:PCI

[88] Subhamoy Maitra, C. E. Veni Madhavan, and Ramarathnam Venkatesan, editors. Progress in Cryptology — INDOCRYPT 2005: 6th International Conference on Cryptology in India, Bangalore, India, December 10– 12, 2005. Proceedings, volume 3797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30805-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3797.

Bebis:2006:AVC

[89] George Bebis, Richard Boyle, Darko Koracin, and Bahram Parvin, editors. Advances in visual computing: first international symposium, ISVC 2005, Lake Tahoe, NV, USA, December 5– 7, 2005. proceedings, volume 3804 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30750-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3804.

Bussler:2006:TSI

[90] Christoph Bussler and Ming-Chien Shan, editors. Technologies for e-services: 6th international workshop, TES 2005, Trondheim, Norway, September 2–3, 2005: revised selected papers, volume 3811 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31067-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .T43 2005. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3811.

Bussler:2006:BPM

[91] Christoph Bussler and Armin Haller, editors. Business Process Management Workshops: BPM 2005 International Workshops, BPI, BPD, ENEI, BPRM, WSCOBPM, BPS, Nancy, France, September 5, 2005. Revised Selected Papers, volume 3812 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32595-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3812.

Faundez-Zanuy:2006:NAA

[92] Marcos Faundez-Zanuy, Léonard Janer, Anna Esposito, Antonio Satue-Villar, Josep Roure, and Virginia Espinosa-Duro, editors. Nonlinear Analyses and Algorithms for Speech Processing: International Conference on Non-Linear Speech Processing, NOLISP 2005, Barcelona, Spain, April 19–22, 2005, Revised Selected Papers, volume 3817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31257-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3817.

Hentenryck:2006:PAD

[93] Pascal Hentenryck, editor. Practical Aspects of Declarative Languages: 8th International Symposium, PADL 2006, Charleston, SC, USA, January 9–10, 2006. Proceedings, volume 3819 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30947-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3819.

Ramanujam:2006:FFS

[94] R. Ramanujam and Sandeep Sen, editors. FSTTCS 2005: Foundations of Software Technology and Theoretical Computer Science: 25th International Conference, Hyderabad, India, December 15–18, 2005. Proceedings, volume 3821 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30495-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3821.

Feng:2006:ISC

[95] Dengguo Feng, Dongdai Lin, and Moti Yung, editors. Information Security and Cryptology: First SKLOIS Conference, CISC 2005, Beijing, China, December 15–17, 2005. Proceedings, volume 3822 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30855-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3822.

Benatallah:2006:SOC

[96] Boualem Benatallah, Fabio Casati, and Paolo Traverso, editors. Service-oriented computing — ICSOC 2005: third international conference, Amsterdam, The Netherlands, December 12–15, 2005, proceedings, volume 3826 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30817-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3826.

Deng:2006:ACI

[97] Xiaotie Deng and Dingzhu Du, editors. Algorithms and computation: 16th international symposium, ISAAC 2005, Sanya, Hainan, China, December 19– 21, 2005, proceedings, volume 3827 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30935-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3827.

Deng:2006:INE

[98] Xiaotie Deng and Yinyu Ye, editors. Internet and network economics: First International workshop, Wine 2005, Hong Kong, China, December 15–17, 2005. proceedings, volume 3828 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30900-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3828.

Weyns:2006:EMA

[99] Danny Weyns, H. VanDyke Parunak, and Fabien Michel, editors. Environments for Multi-Agent Systems II: Second International Workshop, E4MAS 2005, Utrecht, The Netherlands, July 25, 2005, Selected Revised and Invited Papers, volume 3830 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32614-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3830.

Wiedermann:2006:STP

[100] Jiří Wiedermann, Gerard Tel, Jaroslav Pokorný, Mária Bieliková, and Július Štuller, editors. SOFSEM 2006: Theory and Practice of Computer Science: 32nd Conference on Current Trends in Theory and Practice of Computer Science, Merin, Czech Republic, January 21–27, 2006. Proceedings, volume 3831 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31198-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3831.

Zhang:2006:ABH

[101] David Zhang and Anil K. Jain, editors. Advances in biometrics: international conference, ICB 2006, Hong Kong, China, January 5–7, 2006: proceedings, volume 3832 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31111-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.B56 I57 2006. URL http:

//www.springerlink.com/openurl.

                                        asp?genre=issue&issn=0302-9743&
                                        volume=3832.

Filliatre:2006:TPP

[102] Jean-Christophe Filliâtre, Christine Paulin-Mohring, and Benjamin Werner, editors. Types for proofs and programs: international workshop, TYPES 2004, Jouy-en-Josas, France, December 15– 18, 2004: revised selected papers, volume 3839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31428-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A96 T96 2004. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-9743&
                                        volume=3839.

Zhou:2006:FWR

[103] Xiaofang Zhou, Jianzhong Li, Heng Tao Shen, Masaru Kitsuregawa, and Yanchun Zhang, editors. Frontiers of WWW research and development–APWeb 2006: 8th Asia-Pacific Web Conference, Harbin, China, January 16–18, 2006: proceedings, volume 3841 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31142-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .A785 2006. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-9743&
                                        volume=3841.

Shen:2006:AWN

[104] Heng Tao Shen, Jinbao Li, Minglu Li, Jun Ni, and Wei Wang, editors. Advanced Web and network technologies, and applications: APWeb 2006 international workshops: XRA, IWSN, MEGA, and ICSE, Harbin, China, January 16–18, 2006: proceedings, volume 3842 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31158-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.888 .A7852 2006. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3842.

Healy:2006:GDI

[105] Patrick Healy and Nikola S. Nikolov, editors. Graph drawing: 13th international symposium, GD 2005, Limerick, Ireland, September 12–14, 2005, revised papers, volume 3843 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31425-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3843.

Bruel:2006:SEM

[106] Jean-Michel Bruel, editor. Satellite Events at the MoDELS 2005 Conference: MoDELS 2005 International Workshops OCLWS, MoDeVA, MARTES, AOM, MTiP, WiSME, MODAUI, NfC, MDD, WUsCAM, Montego Bay, Jamaica, October 2– 7, 2005, Revised Selected Papers, volume 3844 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-31780-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.

                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3844.

Farre:2006:IAA

[107] Jacques Farré, Igor Litovsky, and Sylvain Schmitz, editors. Implementation and application of automata: 10th international conference, CIAA 2005, Sophia Antipolis, France, June 27–29, 2005, revised selected papers, volume 3845 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31023-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3845.

Herik:2006:CGI

[108] H. Jaap Herik, Yngvi Björnsson, and Nathan S. Netanyahu, editors. Computers and Games: 4th International Conference, CG 2004, Ramat-Gan, Israel, July 5–7, 2004. Revised Papers, volume 3846 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32488-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3846.

Jantke:2006:FWI

[109] Klaus P. Jantke, Aran Lunzer, Nicolas Spyratos, and Yuzuru Tanaka, editors. Federation over the Web: International Workshop, Dagstuhl Castle, Germany, May 1–6, 2005. Revised Selected Papers, volume 3847 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31018-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3847.

Boulicaut:2006:CBM

[110] Jean-François Boulicaut, Luc Raedt, and Heikki Mannila, editors. Constraint-Based Mining and Inductive Databases: European Workshop on Inductive Databases and Constraint Based Mining, Hinterzarten, Germany, March 11– 13, 2004, Revised Selected Papers, volume 3848 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31331-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3848.

Bloch:2006:FLA

[111] Isabelle Bloch, Alfredo Petrosino, and Andrea G. B. Tettamanzi, editors. Fuzzy Logic and Applications: 6th International Workshop, WILF 2005, Crema, Italy, September 15–17, 2005, Revised Selected Papers, volume 3849 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32529-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3849.

Freund:2006:MCI

[112] Rudolf Freund, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa, editors. Membrane Computing: 6th International Workshop, WMC 2005, Vienna, Austria, July 18–21, 2005, Revised Selected and Invited Papers, volume 3850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-30948-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3850.

Narayanan:2006:CVAa

[113] P. J. Narayanan, Shree K. Nayar, and Heung-Yeung Shum, editors. Computer Vision — ACCV 2006: 7th Asian Conference on Computer Vision, Hyderabad, India, January 13–16, 2006, Proceedings, Part I, volume 3851 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31219-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3851.

Narayanan:2006:CVAb

[114] P. J. Narayanan, Shree K. Nayar, and Heung-Yeung Shum, editors. Computer Vision — ACCV 2006: 7th Asian Conference on Computer Vision, Hyderabad, India, January 13–16, 2006, Proceedings, Part II, volume 3852 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31244-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3852.

Ijspeert:2006:BIA

[115] Auke Jan Ijspeert, Toshimitsu Masuzawa, and Shinji Kusumoto, editors. Biologically Inspired Approaches to Advanced Information Technology: Second International Workshop, BIOA-DIT 2006, Osaka, Japan, January 26–27, 2006, volume 3853 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31253-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3853.

Stavrakakis:2006:ACS

[116] Ioannis Stavrakakis and Michael Smirnov, editors. Autonomic communication: Second international IFIP workshop, WAC 2005, Athens, Greece, October 2–5, 2005, revised selected papers, volume 3854 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32992-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3854.

Namjoshi:2006:VMC

[117] Kedar S. Namjoshi E. Allen Emerson, editor. Verification, Model Checking, and Abstract Interpretation: 7th International Conference, VMCAI 2006, Charleston, SC, USA, January 8–10, 2006. Proceedings, volume 3855 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31139-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3855.

Fossorier:2006:AAA

[118] Marc Fossorier, Hideki Imai, Shu Lin, and Alain Poli, editors. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes: 16th International Symposium, AAECC-16, Las Vegas, NV, USA, February 20–24, 2006. Proceedings, volume 3857 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31423-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3857.

Valdes:2006:RAI

[119] Alfonso Valdes and Diego Zamboni, editors. Recent advances in intrusion detection: 8th international symposium, RAID 2005, Seattle, WA, USA, September 7–9, 2005, revised papers, volume 3858 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31779-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3858.

Pointcheval:2006:TCC

[120] David Pointcheval, editor. Topics in Cryptology — CT-RSA 2006: The Cryptographers’ Track at the RSA Conference 2006, San Jose, CA, USA, February 13–17, 2006, proceedings, volume 3860 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31033-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3860.

Dix:2006:FIK

[121] Jürgen Dix and Stephen J. Hegner, editors. Foundations of information and knowledge systems: 4th international symposium, FoIKS 2006, Budapest, Hungary, February 14–17, 2006: proceedings, volume 3861 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31782-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .F65 2006. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3861.

Bordini:2006:PMA

[122] Rafael H. Bordini, Mehdi Dastani, Jürgen Dix, and Amal ElFallah Seghrouchni, editors. Programming Multi-Agent Systems: Third International Workshop, ProMAS 2005, Utrecht, The Netherlands, July 26, 2005, Revised and Invited Papers, volume 3862 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32616-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3862.

Kohlhase:2006:MKM

[123] Michael Kohlhase, editor. Mathematical knowledge management: 4th international conference, MKM 2005, Bremen, Germany, July 15–17, 2005: revised selected papers, volume 3863 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31430-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .M56 2005; QA76.95 .I565 2005. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-9743&
                                        volume=3863.

Shen:2006:CSC

[124] Weiming Shen, Kuo-Ming Chao, Zongkai Lin, Jean-Paul A. Barthès, and Anne James, editors. Computer Supported Cooperative Work in Design II: 9th International Conference, CSCWD 2005, Coventry, UK, May 24–26, 2005, Revised Selected Papers, volume 3865 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32969-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3865.

Dimitrakos:2006:FAS

[125] Theo Dimitrakos, Fabio Martinelli, Peter Y. A. Ryan, and Steve Schneider, editors. Formal Aspects in Security and Trust: Third International Workshop, FAST 2005, Newcastle upon Tyne, UK, July 18–19, 2005, Revised Selected Papers, volume 3866 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32628-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3866.

Romer:2006:WSN

[126] Kay Römer, Holger Karl, and Friedemann Mattern, editors. Wireless sensor networks: third European workshop, EWSN 2006, Zürich, Switzerland, February 13–15, 2006: proceedings, volume 3868 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32158-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7872.D48 E94 2006. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3868.

Bengio:2006:MLM

[127] Samy Bengio and Steve Renals, editors.

Machine learning for multimodal interaction: second international workshop, MLMI 2005, Edinburgh, UK, July 11– 13, 2005: revised selected papers, volume 3869 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32549-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .W69 2005. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3869.

Spaccapietra:2006:JDS

[128] Stefano Spaccapietra, Paolo Atzeni, Wesley W. Chu, Tiziana Catarci, and Katia P. Sycara, editors. Journal on Data Semantics. V, volume 3870 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-31426-1. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 236 pp. LCCN QA76.7 .J68 2006. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3870.

Talbi:2006:AEI

[129] El-Ghazali Talbi, Pierre Liardet, and Pierre Collet, editors. Artificial evolution: 7th international conference, evolution artificielle, EA 2005, Lille, France, October 26–28, 2005, revised selected papers, volume 3871 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33590-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3871.

Bunke:2006:DAS

[130] Horst Bunke and A. Lawrence Spitz, editors. Document analysis systems VII: 7th international workshop, DAS 2006, Nelson, New Zealand, February 13–15, 2006: proceedings, volume 3872 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32140-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7882.P3 I55 2006; TA1650 .D37 2006. URL http:

//www.springerlink.com/openurl.

asp?genre=issue&issn=0302-9743&
volume=3872.

Maicher:2006:CTM

[131] Lutz Maicher and Jack Park, editors.

Charting the Topic Maps Research and Applications Landscape: First International Workshop on Topic Map Research and Applications, TMRA 2005, Leipzig, Germany, October 6–7, 2005, Revised Selected Papers, volume 3873 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32527-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3873.

Missaoui:2006:FCA

[132] Rokia Missaoui and Jürg Schmidt, editors. Formal Concept Analysis: 4th International Conference, ICFCA 2006, Dresden, Germany, February 13– 17, 2006. Proceedings, volume 3874 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32203-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3874.

Ur:2006:HSV

[133] Shmuel Ur, Eyal Bin, and Yaron Wolfsthal, editors. Hardware and Software, Verification and Testing: First International Haifa Verification Conference, Haifa, Israel, November 13–16, 2005, Revised Selected Papers, volume 3875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32604-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3875.

Halevi:2006:TCT

[134] Shai Halevi and Tal Rabin, editors. Theory of Cryptography: Third Theory of Cryptography Conference, TCC 2006, New York, NY, USA, March 4– 7, 2006. Proceedings, volume 3876 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32731-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3876.

Detyniecki:2006:AMR

[135] Marcin Detyniecki, Joemon M. Jose, Andreas Nürnberger, and C. J. ‘Keith’ Rijsbergen. Adaptive Multimedia Retrieval: User, Context, and Feedback: Third International Workshop, AMR 2005, Glasgow, UK, July 28– 29, 2005, Revised Selected Papers, volume 3877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32174-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3877.

Gelbukh:2006:CLI

[136] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 7th International Conference, CICLing 2006, Mexico City, Mexico, February 19–25, 2006. Proceedings, volume 3878 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32205-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3878.

Erlebach:2006:AOA

[137] Thomas Erlebach and Giuseppe Persinao, editors. Approximation and Online Algorithms: Third International Workshop, WAOA 2005, Palma de Mallorca, Spain, October 6–7, 2005, Revised Selected Papers, volume 3879 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32207-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3879.

Rashid:2006:TAO

[138] Awais Rashid and Mehmet Aksit.

Transactions on aspect-oriented software development I, volume 3880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32972-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3880.

Gibet:2006:GHC

[139] Sylvie Gibet, Nicolas Courty, and Jean-François Kamp, editors. Gesture in Human-Computer Interaction and Simulation: 6th International Gesture Workshop, GW 2005, Berder Island, France, May 18–20, 2005, Revised Selected Papers, volume 3881 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32624-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3881.

Lee:2006:DSA

[140] Mong Li Lee, Kian Lee Tan, and Vilas Wuwongse, editors. Database Systems for Advanced Applications: 11th International Conference, DASFAA 2006, Singapore, April 12–15, 2006. Proceedings, volume 3882 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33337-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3882.

Cesana:2006:WSN

[141] atteo Cesana and Luigi Fratta, editors. Wireless Systems and Network Architectures in Next Generation Internet: Second International Workshop of the EURO-NGI Network of Excellence, Villa Vigoni, Italy, July 13–15, 2005, revised selected papers, volume 3883 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34025-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.

com/openurl.asp?genre=issue&issn=
0302-9743&volume=3883.

Durand:2006:SAS

[142] Bruno Durand and Wolfgang Thomas, editors. STACS 2006: 23rd Annual Symposium on Theoretical Aspects of Computer Science, Marseille, France, February 23–25, 2006. Proceedings, volume 3884 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32301-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3884.

Torra:2006:MDA

[143] Vicenç Torra, Yasuo Narukawa, Aïda Valls, and Josep Domingo-Ferrer, editors. Modeling Decisions for Artificial Intelligence: Third International Conference, MDAI 2006, Tarragona, Spain, April 3–5, 2006. Proceedings, volume 3885 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32780-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3885.

Bremer:2006:KDL

[144] Eric G. Bremer, Jörg Hakenberg, Eui-Hong (Sam) Han, Daniel Berrar, and Werner Dubitzky, editors. Knowledge Discovery in Life Science Literature: PAKDD 2006 International Workshop, KDLL 2006, Singapore, April 9, 2006. Proceedings, volume 3886 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32809-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3886.

Correa:2006:LTI

[145] José R. Correa, Alejandro Hevia, and Marcos Kiwi, editors. LATIN 2006: Theoretical Informatics: 7th Latin American Symposium, Valdivia, Chile, March 20–24, 2006. Proceedings, volume 3887 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32755-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3887.

Draheim:2006:TEA

[146] Dirk Draheim and Gerald Weber, editors. Trends in Enterprise Application Architecture: VLDB Workshop, TEAA 2005, Trondheim, Norway, August 28, 2005, Revised Selected Papers, volume 3888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32734-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3888.

Rosca:2006:ICA

[147] Justinian Rosca, Deniz Erdogmus, José C. Príncipe, and Simon Haykin, editors. Independent Component Analysis and Blind Signal Separation: 6th International Conference, ICA 2006, Charleston, SC, USA, March 5–8, 2006. Proceedings, volume 3889 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32630-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3889.

Thompson:2006:DAM

[148] Simon G. Thompson and Robert Ghanea-Hercock, editors. Defence Applications of Multi-Agent Systems: International Workshop, DAMAS 2005, Utrecht, The Netherlands, July 25, 2005, Revised and Invited Papers, volume 3890 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32832-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3890.

Sichman:2006:MAB

[149] Jaime S. Sichman and Luís Antunes, editors. Multi-Agent-Based Simulation VI: International Workshop, MABS 2005, Utrecht, The Netherlands, July 25, 2005, Revised and Invited Papers, volume 3891 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33380-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3891.

Carbone:2006:DCI

[150] Alessandra Carbone and Niles A. Pierce, editors. DNA computing: 11th international workshop on DNA computing, DNA11, London, on, Canada, June 6–9, 2005. revised selected papers, volume 3892 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34161-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3892.

Atzori:2006:VCP

[151] Luigi Atzori, Daniel D. Giusto, and Riccardo Leonardi, editors. Visual content processing and representation: 9th international workshop, VLBV 2005, Sardinia, Italy, September 15–16, 2005, revised selected papers, volume 3893 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33578-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3893.

Grass:2006:ACS

[152] Werner Grass, Bernhard Sick, and Klaus Waldschmidt, editors. Architecture of Computing Systems — ARCS 2006: 19th International Conference, Frankfurt/Main, Germany, March 13– 16, 2006. Proceedings, volume 3894 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32765-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3894.

Goldreich:2006:TCS

[153] Oded Goldreich, Arnold L. Rosenberg, and Alan L. Selman, editors. Theoretical Computer Science: Essays in Memory of Shimon Even, volume 3895 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32880-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3895.

Ioannidis:2006:ADT

[154] Yannis Ioannidis, Marc H. Scholl, Joachim W. Schmidt, Florian Matthes, Mike Hatzopoulos, Klemens Boehm, Alfons Kemper, Torsten Grust, and Christian Boehm, editors. Advances in Database Technology — EDBT 2006: 10th International Conference on Extending Database Technology, Munich, Germany, 26–31 March, 2006, volume 3896 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32960-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3896.

Preneel:2006:SAC

[155] Bart Preneel and Stafford Tavares, editors. Selected Areas in Cryptography: 12th International Workshop, SAC 2005, Kingston, ON, Canada, August 11–12, 2005, Revised Selected Papers, volume 3897 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33108-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3897.

Tuyls:2006:LAM

[156] Karl Tuyls, Pieter Jan ’t Hoen, Katja Verbeeck, and Sandip Sen, editors. Learning and Adaption in Multi-Agent Systems: First International Workshop, LAMAS 2005, Utrecht, The Netherlands, July 25, 2005, Revised Selected Papers, volume 3898 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33053-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3898.

Frintrup:2006:VVA

[157] Simone Frintrup, editor. VOCUS: A Visual Attention System for Object Detection and Goal-Directed Search, volume 3899 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-32759-2. ISSN 0302-9743 (print), 1611-3349 (electronic). 199 (est.) pp. LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3899.

Toni:2006:CLM

[158] Francesca Toni and Paolo Torroni, editors. Computational Logic in Multi-Agent Systems: 6th International Workshop, CLIMA VI, London, UK, June 27–29, 2005: Revised Selected and Invited Papers, volume 3900 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33996-5 (paperback). LCCN QA76.9.L63 C67 2005.

Hill:2006:LBP

[159] Patricia M. Hill, editor. Logic Based Program Synthesis and Transformation: 15th International Symposium, LOP-STR 2005, London, UK, September 7–9, 2005, Revised Selected Papers, volume 3901 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-32654-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3901.

Kronland-Martinet:2006:CMM

[160] Richard Kronland-Martinet, Thierry Voinier, and Sølvi Ystad, editors. Computer Music Modeling and Retrieval: Third International Symposium, CMMR 2005, Pisa, Italy, September 26–28, 2005. Revised Papers, volume 3902 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34027-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3902.

Chen:2006:ISP

[161] Kefei Chen, Robert Deng, Xuejia Lai, and Jianying Zhou, editors. Information Security Practice and Experience: Second International Conference, ISPEC 2006, Hangzhou, China, April 11–14, 2006. Proceedings, volume 3903 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33052-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3903.

Baldoni:2006:DAL

[162] Matteo Baldoni, Ulle Endriss, Andrea Omicini, and Paolo Torroni, editors. Declarative Agent Languages and Technologies III: Third International Workshop, DALT 2005, Utrecht, The Netherlands, July 25, 2005, Selected and Revised Papers, volume 3904 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33106-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3904.

Collet:2006:GPE

[163] Pierre Collet, Marco Tomassini, Marc Ebner, Steven Gustafson, and Anikó Ekárt, editors. Genetic Programming: 9th European Conference, EuroGP 2006, Budapest, Hungary, April 10–12, 2006. Proceedings, volume 3905 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33143-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.

com/openurl.asp?genre=issue&issn=
0302-9743&volume=3905.

Gottlieb:2006:ECC

[164] Jens Gottlieb and Günther R. Raidl, editors. Evolutionary Computation in Combinatorial Optimization: 6th European Conference, EvoCOP 2006, Budapest, Hungary, April 10– 12, 2006. Proceedings, volume 3906 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33178-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3906.

Rothlauf:2006:AEC

[165] Franz Rothlauf, Jürgen Branke, Stefano Cagnoni, Ernesto Costa, Carlos Cotta, Rolf Drechsler, Evelyne Lutton, Penousal Machado, Jason H. Moore, Juan Romero, George D. Smith, Giovanni Squillero, and Hideyuki Takagi, editors. Applications of Evolutionary Computing: EvoWorkshops 2006: EvoBIO, EvoCOMNET, EvoHOT, EvoIASP, EvoINTERACTION, EvoMUSART, and EvoSTOC, Budapest, Hungary, April 10–12, 2006. Proceedings, volume 3907 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33237-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3907.

Bui:2006:IIC

[166] Alain Bui, Marc Bui, and Thomas Böhme, editors. Innovative Internet Community Systems: 5th International Workshop, IICS 2005, Paris, France, June 20–22, 2005. Revised Papers, volume 3908 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33973-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3908.

Apostolico:2006:RCM

[167] Alberto Apostolico, Concettina Guerra, Sorin Istrail, Pavel Pevzner, and Michael Waterman, editors. Research in Computational Molecular Biology: 10th Annual International Conference, RE-COMB 2006, Venice, Italy, April 2– 5, 2006. Proceedings, volume 3909 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33295-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3909.

Wyrzykowski:2006:PPA

[168] Roman Wyrzykowski, Jack Dongarra, and Norbert Meyer, editors. Parallel Processing and Applied Mathematics: 6th International Conference, PPAM 2005, Poznan, Poland, September 11–14, 2005, Revised Selected Papers, volume 3911 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34141-2 (soft cover). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3911.

Garcia:2006:SEM

[169] Alessandro Garcia, Ricardo Choren, Carlos Lucena, Paolo Giorgini, Tom Holvoet, and Alexander Romanovsky, editors. Software Engineering for Multi-Agent Systems IV: Research Issues and Practical Applications, volume 3914 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33580-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 255 pp. LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3914.

Nayak:2006:KDX

[170] Richi Nayak and Mohammed J. Zaki, editors. Knowledge Discovery from XML Documents: First International Workshop, KDXD 2006, Singapore, April 9, 2006. Proceedings, volume 3915 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33180-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3915.

Li:2006:DMB

[171] Jinyan Li, Qiang Yang, and Ah-Hwee Tan, editors. Data Mining for Biomedical Applications: PAKDD 2006 Workshop, BioDM 2006, Singapore, April 9, 2006. Proceedings, volume 3916 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33104-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3916.

Chen:2006:ISI

[172] Hsinchun Chen, Fei Yue Wang, Christopher C. Yang, Daniel Zeng, Michael Chau, and Kuiyu Chang, editors. Intelligence and Security Informatics: International Workshop, WISI 2006, Singapore, April 9, 2006. Proceedings, volume 3917 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33361-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3917.

Ng:2006:AKD

[173] Wee Keong Ng, Masaru Kitsuregawa, Jianzhong Li, and Kuiyu Chang, editors. Advances in Knowledge Discovery and Data Mining: 10th Pacific–Asia Conference, PAKDD 2006, Singapore, April 9–12, 2006. Proceedings, volume 3918 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33206-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3918.

Hermanns:2006:TAC

[174] Holger Hermanns and Jens Palsberg, editors. Tools and Algorithms for the Construction and Analysis of Systems: 12th International Conference, TACAS 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 25– April 2, 2006. Proceedings, volume 3920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33056-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3920.

Aceto:2006:FSS

[175] Luca Aceto and Anna Ingólfsdóttir, editors. Foundations of Software Science and Computation Structures: 9th International Conference, FOSSACS 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 25–31, 2006. Proceedings, volume 3921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33045-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3921.

Baresi:2006:FAS

[176] Luciano Baresi and Reiko Heckel, editors. Fundamental Approaches to Software Engineering: 9th International Conference, FASE 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 27–28, 2006. Proceedings, volume 3922 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33093-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.

                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3922.

Mycroft:2006:CCI

[177] Alan Mycroft and Andreas Zeller, editors. Compiler Construction: 15th International Conference, CC 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 30–31, 2006. Proceedings, volume 3923 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33050-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3923.

Sestoft:2006:PLS

[178] Peter Sestoft, editor. Programming Languages and Systems: 15th European Symposium on Programming, ESOP 2006, Held as Part of the Joint European Conferences on Theory and Practice of Software, ETAPS 2006, Vienna, Austria, March 27– 28, 2006. Proceedings, volume 3924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33095-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3924.

Valmari:2006:MCS

[179] Antti Valmari, editor. Model Checking Software: 13th International SPIN Workshop, Vienna, Austria, March 30– April 1, 2006. Proceedings, volume 3925 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33102-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3925.

Hespanha:2006:HSC

[180] João Hespanha and Ashish Tiwari, editors. Hybrid Systems: Computation and Control: 9th International Workshop, HSCC 2006, Santa Barbara, CA, USA, March 29–31, 2006. Proceedings, volume 3927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33170-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3927.

Domingo-Ferrer:2006:SCR

[181] Josep Domingo-Ferrer, Joachim Posegga, and Daniel Schreckling, editors. Smart Card Research and Advanced Applications: 7th IFIP WG 8.8/11.2 International Conference, CARDIS 2006, Tarragona, Spain, April 19– 21, 2006. Proceedings, volume 3928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33311-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3928.

MacCaull:2006:RMC

[182] Wendy MacCaull, Michael Winter, and Ivo Düntsch, editors. Relational Methods in Computer Science: 8th International Seminar on Relational Methods in Computer Science, 3rd International Workshop on Applications of Kleene Algebra, and Workshop of Cost Action 274: Tarski, St. Catharines, ON, Canada, February 22–26, 2005, Selected Revised Papers, volume 3929 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33339-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3929.

Apolloni:2006:NN

[183] Bruno Apolloni, Maria Marinaro, Giuseppe Nicosia, and Roberto Tagliaferri, editors. Neural Nets: 16th Italian Workshop on Neural Nets, WIRN 2005, and International Workshop on Natural and Artificial Immune Systems, NAIS 2005, Vietri sul Mare, Italy, June 8–11, 2005, Revised Selected Papers, volume 3931 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33183-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3931.

Bonchi:2006:KDI

[184] Francesco Bonchi and Jean-Francois Boulicaut, editors. Knowledge Discovery in Inductive Databases: 4th International Workshop, KDID 2005, Porto, Portugal, October 3, 2005, Revised Selected and Invited Papers, volume 3933 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9.

ISBN 3-540-33292-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3933.

Clark:2006:SPC

[185] John A. Clark, Richard F. Paige, Fiona A. C. Polack, and Phillip J. Brooke, editors. Security in Pervasive Computing: Third International Conference, SPC 2006, York, UK, April 18– 21, 2006. Proceedings, volume 3934 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33376-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3934.

Won:2006:ISC

[186] Dongho Won and Seungjoo Kim, editors. Information Security and Cryptology — ICISC 2005: 8th International Conference, Seoul, Korea, December 1–2, 2005, Revised Selected Papers, volume 3935 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33354-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3935.

Lalmas:2006:AIR

[187] Mounia Lalmas, Andy MacFarlane, Stefan Rüger, Anastasios Tombros, Theodora Tsikrika, and Alexei Yavlinsky, editors. Advances in Information Retrieval: 28th European Conference on IR Research, ECIR 2006, London, UK, April 10–12, 2006. Proceedings, volume 3936 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33347-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3936.

Priami:2006:TCS

[188] Corrado Priami, Luca Cardelli, and Stephen Emmott, editors. Transactions on Computational Systems Biology IV, volume 3939 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33245-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3939.

Saunders:2006:SLS

[189] Craig Saunders, Marko Grobelnik, and Steve Gunn, editors. Subspace, Latent Structure and Feature Selection: Statistical and Optimization Perspectives Workshop, SLSPS 2005, Bohinj, Slovenia, February 23–25, 2005, Revised Selected Papers, volume 3940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34137-4 (softcover). LCCN ????

Gilroy:2006:ISD

[190] Stephen W. Gilroy and Michael D. Harrison, editors. Interactive Systems: Design, Specification, and Verification. 12th International Workshop, DSV-IS 2005, Newcastle upon Tyne, UK, July 13–15, 2005. Revised Papers, volume 3941 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34145-5 (softcover). LCCN ????

Pan:2006:TLD

[191] Zhigeng Pan, Ruth Aylett, Holger Diener, Xiaogang Jin, Stefan Göbel, and Li Li, editors. Technologies for E-Learning and Digital Entertainment: First International Conference, Edutainment 2006, Hangzhou, China, April 16–19, 2006. Proceedings, volume 3942 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33423-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3942.

Guelfi:2006:RIS

[192] Nicolas Guelfi and Anthony Savidis, editors. Rapid Integration of Software Engineering Techniques: Second International Workshop, RISE 2005, Heraklion, Crete, Greece, September 8–9, 2005, Revised Selected Papers, volume 3943 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-34063-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3943.

Hagiya:2006:FLP

[193] Masami Hagiya and Philip Wadler, editors. Functional and Logic Programming: 8th International Symposium, FLOPS 2006, Fuji-Susono, Japan, April 24–26, 2006. Proceedings, volume 3945 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-33438-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

                                        ???? URL http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3945.

Roth-Berghofer:2006:MRC

[194] Thomas R. Roth-Berghofer, Stefan Schulz, and David B. Leake, editors. Modeling and Retrieval of Context: Second International Workshop, MRC 2005, Edinburgh, UK, July 31–August 1, 2005, Revised Selected Papers, volume 3946 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33587-0 (softcover). LCCN QA765.758 .I577 2005.

Chung:2006:AGP

[195] Yeh-Ching Chung and José E. Moreira, editors. Advances in Grid and Pervasive Computing: First International Conference, GPC 2006, Taichung, Taiwan, May 3–5, 2006: Proceedings, volume 3947 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33809-8 (paperback). LCCN QA76.9.C58 G63 2006.

Barthe:2006:CAS

[196] Gilles Barthe, Benjamin Gregoire, and Marieke Huisman, editors. Construction and analysis of Safe, Secure, and Interoperable Smart Devices: Second International Workshop, CASSIS 2005, Nice, France, March 8–11, 2005, Revised Selected Papers, volume 3956 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33689-3 (softcover). LCCN ????

Yung:2006:PKC

[197] Moti Yung, Yevgeniy Dodis, Aggelos Kiayias, and Tal Malkin, editors. Public Key Cryptography: PPK 2006: 9th International Conference on Theory and Practice in Public-Key Cryptography, New York, NY, USA, April 24–26, 2006. Proceedings, volume 3958 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33851-9 (softcover). LCCN

???? URL http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=3958.

Cai:2006:TAM

[198] Jin-Yi Cai, S. Barry Cooper, and Angsheng Li, editors. Theory and Applications of Models of Computation: Third International Conference, TAMC 2006, Beijing, China, May 15–20, 2006, Proceedings, volume 3959 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34021-1 (softcover). LCCN ????

Vieira:2006:CPP

[199] Renata Vieira et al., editors. Computational Processing of the Portuguese Language: 7th International Workshop, PROPOR 2006, Itatiaia, Brazil, May 13–17, 2006: Proceedings, volume 3960 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34045-9 (paperback). LCCN PC5044.5 .I58 2006.

IJsselsteijn:2006:PTF

[200] Wijnand IJsselsteijn, Yvonne de Kort, and Cees Midden, editors. Persuasive Technology: First International Conference on Persuasive Technology for Human Well-Being, PERSUASIVE 2006, Eindhoven, The Netherlands, May 2006, Proceedings, volume 3962 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34291-5 (softcover). LCCN ????

Uyar:2006:TCS

[201] M. Uemit Uyar, Ali Y. Duale, and Mariusz A. Fecko, editors. Testing of Communicating Systems: 18th IFIP TC 6/WG 6.1 International Conference, TESTCOM 2006, New York, NY, USA, May 16–18, 2006, Proceedings, volume 3964 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34184-6 (softcover). LCCN ????

Cimatti:2006:FMH

[202] Alessandro Cimatti and Marco Bernardo, editors. Formal methods for hardware verification: 6th International School on Formal Methods for the Design of Computer, Communication, and Software Systems, SFM 2006, Bertinoro, Italy, May 22–27, 2006: advanced lectures, volume 3965 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34304-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.F67 I586 2006. URL http:/

/www.loc.gov/catdir/enhancements/
fy0661/2006925529-d.html;   http:
//www.loc.gov/catdir/toc/fy0705/
2006925529.html.

Grigoriev:2006:CST

[203] Dima Grigoriev, John Harrison, and Edward A. Hirsch, editors. Computer Science — Theory and Applications: First International Symposium on Computer Science in Russia, Csr 2006, St. Petersburg, Russia, June 8–12. 2006, Proceedings, volume 3967 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34166-8 (softcover). LCCN ????

Fishkin:2006:PCI

[204] Kenneth P. Fishkin, Bernt Schiele, and Paddy Nixon, editors. Pervasive Computing: 4th International Conference, PERVASIVE 2006, Dublin, Ireland, May 7–10, 2006: Proceedings, volume 3968 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33894-2 (paperback). LCCN QA76.59 .P47 2006.

Ytrehus:2006:CCI

[205] Øyvind Ytrehus, editor. Coding and Cryptography: International Workshop, WCC 2005, Bergen, Norway, March 14– 18, 2005. Revised Selected Papers, volume 3969 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35481-6. LCCN QA76.9.A25 I557 2005.

Braun:2006:WWI

[206] Torsten Braun, Georg Carle, and Sonia Fahmy, editors. Wired/Wireless Internet Communications: 4th International Conference, WWIC 2006, Bern, Switzerland, May 10–12, 2006, Proceedings, volume 3970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34023-8 (softcover). LCCN ????

Wang:2006:ANN

[207] Jun Wang, Zhang Yi, and Jacek M. Zurada, editors. Advances in Neural Networks: ISNN 2006: Third International Symposium on Neural Networks, ISNN 2006, Chengdu, China, May 28: June 1, 2006, Proceedings, Part I, volume 3971– 3973 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34439-X (part 1), 3-540-34437-3 (part 2), 3-540-34482-9 (part 3). LCCN ????

Mehrotra:2006:ISI

[208] Sharad Mehrotra, Daniel D. Zeng, and Hsinchun Chen, editors. Intelligence and Security Informatics: IEEE International Conference on Intelligence and Security Informatics, ISI 2006, San Diego, CA, USA, May 23–24, 2006. Proceedings, volume 3975 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34478-0 (softcover). LCCN ????

Boavida:2006:NNT

[209] Fernando Boavida, Thomas Plagemann, and Burkhard Stiller, editors. Networking 2006. Networking Technologies, Services, and Protocols; Performance of Computer and Communication Networks; Mobile and Wireless Communications Systems: 5th International IFIP-TC 6 Networking Conference, Coimbra, Portugal, May 15– 19, 2006, Proceedings, volume 3976 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34192-7 (softcover). LCCN ????

Huang:2006:CVH

[210] Thomas S. Huang, Nicu Sebe, and Michael S. Lew, editors. Computer Vision in Human-Computer Interaction: ECCV 2006 Workshop on HCI, Graz, Austria, May 13, 2006. Proceedings, volume 3979 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34202-8 (softcover). LCCN ????

Gavrilova:2006:CSA

[211] Marina Gavrilova, Osvaldo Gervasi, and Vipin Kumar, editors. Pervasive Computing: 4th International Conference, Pervasive 2006, Dublin, Ireland, May 7–10, 2006, Proceedings, volume 3980 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34070-X (softcover). LCCN ????

Hazas:2006:LCA

[212] Mike Hazas, John Krumm, and Thomas Strang, editors. Location- and Context-Awareness: Second International Workshop, LOCA 2006, Dublin, Ireland, May 10–11, 2006, Proceedings, volume 3987 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34150-1 (softcover). LCCN ????

Beckmann:2006:LAC

[213] Arnold Beckmann, Ulrich Berger, Benedikt Löwe, and John V. Tucker, editors. Logical approaches to computational barriers: Second Conference on Computability in Europe, CiE 2006, Swansea, UK, June 30–July 5, 2006: proceedings, volume 3988 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006.

ISBN 3-540-35466-2, 3-540-35468-9 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.59 .C67

                                        2006. URL http://link.springer.
                                        com/book/10.1007/11780342.

Beck:2006:IAT

[214] J. Christopher Beck and Barbara M. Smith, editors. Integration of AI and OR Techniques in Constraint Programming for Combinatorial Optimization Problems: Third International Conference, CPAIOR 2006, Cork, Ireland, May 31– June 2, 2006, Proceedings, volume 3990 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34306-7 (softcover). LCCN ????

Alexandrov:2006:CSI

[215] Vassil N. Alexandrov, Geert Dick van Albada, and Peter M. A. Sloot, editors. Computational Science: ICCS 2006: 6th International Conference, Reading, UK, May 28–31, 2006, Proceedings, Part I, volume 3991 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34379-2 (softcover). LCCN ????

Mueller:2006:ETI

[216] Guenter Mueller, editor. Emerging Trends in Information and Communication Security: International Conference, ETRICS 2006, Freiburg, Germany, June 6–9, 2006. Proceedings, volume 3995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34640-6 (softcover). LCCN ????

Grieskamp:2006:FAS

[217] Wolfgang Grieskamp and Carsten Weise, editors. Formal Approaches to Software Testing: 5th International Workshop, FATES 2005, Edinburgh, UK, July 11, 2005, Revised Selected Papers, volume 3997 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34454-3 (softcover). LCCN ????

Calamoneri:2006:ACI

[218] Tiziana Calamoneri, Irene Finocchi, and Giuseppe F. Italiano, editors. Algorithms and Complexity: 6th Italian Conference, CIAC 2006, Rome, Italy, May 29–31, 2006. Proceedings, volume 3998 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34375-X (softcover). LCCN ????

Stanczak:2006:RAW

[219] Sławomir Stańczak, Marcin Wiczanowski, and Holger Boche, editors. Resource Allocation in Wireless Networks: Theory and Algorithms, volume 4000 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-46248-1 (print), 3-540-46249-X (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). ???? pp. LCCN ????

URL http://www.springerlink.com/
content/978-3-540-46249-1.

Alvarez:2006:EAI

[220] Carme Alvarez and Maria Serna, editors. Experimental algorithms: 5th international workshop, WEA 2006, Cala Galdana, Menorca, Spain, May 24– 27, 2006, proceedings, volume 4007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34597-3 (softcover). LCCN ????

Sure:2006:SWR

[221] York Sure and John Domingue, editors. The Semantic Web: Research and Applications, volume 4011 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34544-2 (softcover). ???? pp. LCCN ????

Eliassen:2006:DAI

[222] Frank Eliassen and Alberto Montresor, editors. Distributed Applications and Interoperable Systems: 6th IFIP WG 6.1 International Conference, DAIS 2006, Athens, Greece, June 14–15, 2006. Proceedings, volume 4025 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35126-4 (softcover). LCCN ????

Gibbons:2006:DCS

[223] Phil Gibbons, Tarek Abdelzaher, and James Aspnes, editors. Distributed Computing in Sensor Systems: Second IEEE International Conference, DCOSS 2006, San Francisco, CA, USA, June 18–20, 2006. Proceedings, volume 4026 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35227-9 (softcover). LCCN ????

Gorrieri:2006:FMO

[224] Roberto Gorrieri and Heike Wehrheim, editors. Formal Methods for Open Object-Based Distributed Systems: 8th IFIP WG 6.1 International Conference,FMOODS 2006, Bologna, Italy, June 14–16, 2006, Proceedings, volume 4037 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34893-X (softcover). LCCN ????

Ciancarini:2006:CML

[225] Paolo Ciancarini and Herbert Wiklicky, editors. Coordination Models and Languages: 8th International Conference, Coordination 2006, Bologna, Italy, June 14–16, 2006. Proceedings, volume 4038 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-34694-5 (softcover). LCCN ????

Atzeni:2006:PKI

[226] Andrea S. Atzeni and Antonio Lioy, editors. Public Key Infrastructure: Third European PKI Workshop: Theory and Practice, EUROPKI 2006, Turin, Italy, June 19–20, 2006. Proceedings, volume 4043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35151-5 (softcover). LCCN ????

Abrahamsson:2006:EPA

[227] Pekka Abrahamsson, Michele Marchesi, and Giancarlo Succi, editors. Extreme programming and Agile Processes in Software Engineering: 7th International Conference, XP 2006, Oulu, Finland, June 17–22, 2006, Proceedings, volume 4044 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-35094-2 (softcover). LCCN ????

Perner:2006:ADM

[228] Petra Perner, editor. Advances in data mining: applications in medicine, web mining, marketing, image and signal mining: 6th Industrial Conference on Data Mining, ICDM 2006, Leipzig, Germany, July 14–15, 2006: proceedings, volume 4065 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-36036-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D343.

Jesshope:2006:ACS

[229] Chris Jesshope and Colin Egan, editors. Advances in computer systems architecture: 11th Asia-Pacific conference, AC-SAC 2006, Shanghai, China, September 6–8, 2006, proceedings, volume 4186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-40056-7 (softcover). LCCN QA76.9.A73 A28 2006.

Min:2006:FHP

[230] Geyong Min et al., editors. Frontiers of high performance computing and networking: ISPA 2006 workshops: ISPA 2006 international workshops, FHPCN, XHPC, S-GRACE, GridGIS, HPC-GTP, PDCE, ParDM-Com, WOMP, ISDF, and UPWN, Sorrento, Italy, December 4-7, 2006: proceedings, volume 4331 of Lecture notes in computer science, 0302-9743. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-49860-5 (paperback). LCCN QA76.58 .I88 2006a; QA76.88 .F76 2006. URL http:/

/www.loc.gov/catdir/enhancements/
fy0825/2006937143-d.html.

Lai:2006:ACA

[231] Xuejia Lai and Kefei Chen, editors. Advances in Cryptology — ASIACRYPT 2006: 12th International Conference on the Theory and Application of Cryptology and Information Security, Shanghai, China, December 3–7, 2006. Proceedings, volume 4284 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CODEN LNCSD9. ISBN 3-540-49475-8 (print), 3-540-49476-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

???? URL http://www.springerlink.
com/content/978-3-540-49476-8.